OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (a1234):

@longboardman15 it is possible to do it that way, but the person should know how to do it without having the choices.

OpenStudy (anonymous):

okay yhanks

OpenStudy (a1234):

Whoops, I'm wrong. I'll see what I did wrong.

OpenStudy (a1234):

I'll use substitution since it is easy to solve for y. 5x + y = 9 This is y = -5x + 9 3x + 2(-5x + 9) = 4 3x -10x + 18 = 4 -7x = -14 x = 2 Put x into x of the second equation. 3(2) + 2y = 4 6 + 2y = 4 2y = -2 y = -1 (2, -1)